Circular No. 5975

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994L IN NGC 2848 On Apr. 13, Nicholas J. Brown, Clarkson, Western Australia, reported

an apparent supernova in NGC 2848 (α = 9^{h}17^{m}.9, δ = −16^{o}19′, equinox 1950.0) that he had discovered visually on Apr. 8.7 UT at m_{v} = 14.7; he stated that the object was located ∼ 30′′ southwest of the galaxy’s center. In response to requests for confirmation by the Central Bureau, L. Pasquini, European Southern Observatory; and M. Osterloh, Max-Planck-Institut für Astronomie, Heidelberg, report: “On Apr. 13.99 UT, we obtained a low-resolution spectrogram (range 370-800 nm, resolution 2 nm) of the supernova candidate in NGC 2848 using the 2.2-m MPI/ESO telescope (+ EFOSC II). The spectrum shows broad Balmer lines with prominent P-Cyg profiles overimposed on a blue continuum. Adopting a redshift velocity of 1795 km/s for the host galaxy, absorption minima are measured at 635, 473, 421, and 382 nm. Additional strong emission lines are also detected at 459 and 496 nm. The object thus appears to be a type-II supernova a few weeks after maximum light.”

COMET TAKAMIZAWA-LEVY (1994f)

Parabolic orbital elements from MPEC 1994-H02:

T = 1994 May 22.602 TT  ω = 61.^{o}791
Ω = 306.903 }2000.0
q = 1.35790 AU  i = 132.993
1994TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Apr. 14 21^{h}23.^{m}37 + 4^{o}15.^{′}2 1.667 1.471 61.^{o}0 36.^{o}6 9.8
16 21 22.65 + 5 40.2 1.615 1.460 62.8 37.7 9.7
18 21 21.78 + 7 10.6 1.562 1.450 64.6 38.7 9.6
20 21 20.73 + 8 47.1 1.510 1.440 66.4 39.7 9.5
22 21 19.49 +10 30.1 1.458 1.430 68.2 40.7 9.4
24 21 18.02 +12 20.5 1.407 1.422 70.0 41.7 9.3
26 21 16.30 +14 18.7 1.356 1.413 71.7 42.5 9.2
28 21 14.28 +16 25.7 1.307 1.405 73.5 43.4 9.1
30 21 11.93 +18 42.1 1.259 1.398 75.3 44.2 9.0
May 2 21 09.18 +21 08.7 1.212 1.391 77.0 44.9 8.9
4 21 05.96 +23 46.2 1.167 1.385 78.7 45.5 8.8
6 21 02.18 +26 35.2 1.125 1.380 80.4 46.1 8.7
8 20 57.74 +29 36.1 1.084 1.375 82.0 46.6 8.6

1994 April 17 Daniel W. E. Green